OpenStudy (anonymous):

Solve for x. x − 6(3 + 4) = 33

OpenStudy (michele_laino):

first we can to do the sum: \(6 \cdot (3+4)=6 \cdot 7=...?\)

OpenStudy (michele_laino):

oops.. we can do the sum

OpenStudy (michele_laino):

please what do you get?

OpenStudy (anonymous):

42

OpenStudy (michele_laino):

that's right, so the equation becomes: \(x-42=33\)

OpenStudy (michele_laino):

next I add to bot sides 42, so I can write this: \[x - 42 + 42 = 33 + 42\] please simplify

OpenStudy (michele_laino):

hint: what is \(-42+42=...?\) and \(33+42=...?\)

OpenStudy (anonymous):

x=75

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

:)